what is x when 100=500-.25x
Show work

Answer by mananth(16946)

100=500-.25x
multiply by 100
10000=50000-25x
add -50000t0 both sides
10000-50000=50000-50000-25x
-40000=-25x
divide by -25
1600=x
..
m.ananth@hotmail.ca
...